OpenStudy (anonymous):

If the ordered pair (-5/2, y) lies on the graph of y = 7 - 2x, find the value of y.

OpenStudy (anonymous):

(-5/2,y) means that x = -5/2 so substitute x = -5/2 and solve for y y = 7 - 2(-5/2) y = 7+5 y = 12

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome.